# Community-level Contagion among Diverse Financial Assets

### v1

@@ Abstract
As global financial markets become increasingly interconnected, financial contagion has developed into a major influencer of asset price dynamics. Motivated by this context, our study explores financial contagion both within and between asset communities. We contribute to the literature by examining the contagion phenomenon at the community level rather than among individual assets. Our experiments rely on high-frequency data comprising cryptocurrencies, stocks and US ETFs over the 4-year period from April 2019 to May 2023. Using the Louvain community detection algorithm, Vector Autoregression contagion detection model and Tracy-Widom random matrix theory for noise removal from financial assets, we present three main findings. Firstly, while the magnitude of contagion remains relatively stable over time, contagion density (the percentage of asset pairs exhibiting contagion within a financial system) increases. This suggests that market uncertainty is better characterized by the transmission of shocks more broadly than by the strength of any single spillover. Secondly, there is no significant difference between intra- and inter-community contagion, indicating that contagion is a system-wide phenomenon rather than being confined to specific asset groups. Lastly, certain communities themselves, especially those dominated by Information Technology assets, consistently appear to act as major contagion transmitters in the financial network over the examined period, spreading shocks with high densities to many other communities. Our findings suggest that traditional risk management strategies such as portfolio diversification through investing in low-correlated assets or different types of investment vehicle might be insufficient due to widespread contagion.
@@
[label2]organization=School of Computing, Dublin City University, addressline=Collins Ave Ext, Whitehall, city=Dublin, postcode=D09 W6Y4, country=Ireland
@@ 1 Introduction
In light of the inherent risks associated with investing, a substantial body of literature has sought to uncover the underlying characteristics of financial markets to help investors in making more informed decisions. These studies focus on enhancing investment strategies by identifying factors that can help in mitigating risks and optimizing returns [11, 12, 13]. One significant area of this topic is the investigation of financial contagion - the phenomenon where economic shocks or market disturbances in one asset, market sector of country’s financial market spread to other assets or sectors in that country or even other countries, leading to price fluctuations in the affected ones and potentially amplifying systemic risk [14]. Understanding these contagion mechanisms is crucial, as they can undermine diversification strategies and exacerbate market volatility. For example, a study conducted by Luchtenberg and Vu [15] discovered the financial contagion effects between 10 countries from North America, Europe and East Asia Pacific during one of the biggest financial crashes in the Global Financial Crisis (GFC) of 2007/08. They revealed that this phenomenon became evident in all considered countries with key economic fundamentals such as trade structure, inflation, interest rates and investor risk aversion being significant determinants of contagion. The results suggest that international diversification by geography alone may not effectively shield investors, and that policymakers in both developed and developing countries must consider contagion risks in their financial strategies. In [16], the same pattern occurred when the contagion has been shown to strike rapidly during the outbreak of the Covid-19 pandemic across global financial markets, especially in developing countries, eventually causing the synchronized downturns of numerous assets across various investment vehicles. To this end, our study contributes to this line of inquiry by examining the contagion dynamics across traditional assets such as stocks and ETFs as well as emerging ones like cryptocurrencies.
@@
This study draws on high-frequency data consisting of 221 time series of closing prices recorded at 30-minute intervals from April 2019 to May 2023. This period is chosen for its inclusion of several major events - both positive and negative - that significantly impact the financial markets. Moreover, it is also a time frame analyzed in related research [12]. The dataset covers 146 stocks, 49 U.S.-based exchange-traded funds (ETFs) tied to corresponding U.S. indices and 27 cryptocurrencies. All of these data were collected from the FirstRate Data platform. Regarding the methodologies, we use Minimum Spanning Tree (MST) and Louvain community detection to identify the community structure of the assets [26, 11]. The contagion signals are detected using the well-known Vector Autoregression (VAR) approach [27]. In addition, measurement noise and inherent noise have been proven to exist in financial markets and potentially impact relevant experimental results [28, 26]. Therefore, to remove this effect from assets, we adopt the Tracy-Widom random matrix theory, which was introduced in [29] and has been widely recognized as an efficient theory for noise removal in various areas, especially financial markets [30, 31, 32].
The remainder of the article is organized as follows: Section 2 presents an overview of related works. Section 3 provides a description of the datasets. Section 4 discusses methodology, preprocessing procedures and metrics. Section 5 describes the experimental results followed by their implications and hypothesis. Lastly, the conclusion of this study is given in Section 6.
@@ 2.1 Upheaval in Financial Markets due to Recent Events
Global financial markets were affected by a series of major events during the 4-year period from April 2019 to May 2023. One of those that generated a great deal of headlines was the US-China trade war, which began in January 2018 with the US imposing tariffs on various Chinese imports, especially in the Information Technology sector [33]. In response, China enacted reciprocal tariffs on the US goods and introduced legislation restricting the use of US-made technology products by Chinese consumers  [34, 35]. The trade war lasted for more than 2 years until January 2020, peaking in intensity during 2019 and negatively affected financial markets in both countries. Sectors such as Commodities, Energy and Technology experienced sharp declines, with major stocks and ETFs suffering substantial losses [36, 37]. Furthermore, as the central hubs of the global economy, the effects of this event spilled over to other parts of the world, increasing the volatility of financial markets worldwide [38].
The next noticeable event is the Covid-19 pandemic [39]. Different studies have explored its impact across various aspects, such as asset correlations, investor behavior, policy responses and market reactions [40, 41, 42]. The same conclusion made by the majority is that the pandemic had an exceptionally strong and widespread effect on global financial markets, leading to one of the most significant economic crises in history. Specifically, this event caused a sharp price decline in a wide range of assets across nearly all types of investment vehicles [40, 41], with most severe effects were observed during the pandemic’s first outbreak, between March and June 2020 [40, 41]. These consequences largely resulted from government interventions aimed at containing the virus, such as international border closures, mobility restrictions, and widespread shutdowns of economic activity [42].
Another event worth mentioning is the ongoing Ukraine-Russia conflict, which occurred on 24 February 2022 [43]. Russia, as major exporter of oil and natural gas, while Ukraine, a key energy transit route and grain supplier, both play important roles in global supply chains, especially for countries in Europe. Consequently, the conflict between the two countries had a significant impact on financial markets, especially in relevant business sectors such as Energy, Financials and Consumer Staples [44, 45]. However, it is shown that the effects were uneven since most severe financial consequences concentrated on European countries due to their strong commercial and political linkages with Russia and Ukraine[46, 47], while the impact on Asian markets and developed economies like the United States was limited or negligible [48, 49, 50]. Notably, the financial impact of this event was found to be modest compared to the past global disruptions like Covid-19 and the US–China trade war [46, 49, 51].
@@ 2.2 Contagion in Financial Markets
The topic of financial contagion has been extensively explored in the literature, with studies examining various dimensions of the phenomenon, such as contagion across countries, market indicators (e.g., interest rates, inflation rates, exchange rates, etc) and individual assets, both within the same investment type and across different types. For instance, Ahmed and Masahiro in [52] examined financial contagion among 9 Asian countries using 3 market indicators: stock market index, interest rate, and exchange rate against the US dollar. They employed the Vector Autoregressive (VAR) model to detect possible contagion effects during the period from July 1997 to June 1998, covering the 1997 Asian Crisis. Their findings revealed contagion in stock indices and currency exchange rates. However, the effect was weak and limited since not all examined countries showed signs of receiving or transmitting contagion while those that did only interacted with a few others. By contrast, they found no evidence of contagion in interest rates. Interestingly, the study noted that while contagion intensified during periods of high market turbulence, it was not the primary driver of the crisis in 1997 due to its weak and limited effect seen among the countries.
@@
As a common method for measuring contagion, Tugba et al. [53] also used the VAR model for their research. In this study, they investigated the shock spillovers between cryptocurrencies, technology stocks, US ETFs and NFTs from March 2018 to September 2021. Their study assessed contagion both within and across different types of investment vehicle. The results showed that contagion primarily occurs within the same investment type since most contagion links are found among assets of the same type. However, these contagions do appear between technology stocks and US ETFs, while very little spillover is found between cryptocurrencies and traditional assets. In addition, NFTs show no contagion with other investment types.
@@
Apart from VAR, different methods have been introduced to examine financial contagion across different time periods and market conditions. For instance, the authors in [19] employed DCC-GARCH (Dynamic Conditional Correlation GARCH) and Wavelet Cross-Correlation methods to examine the contagion across various cryptocurrencies and equity markets in both developing and giant economies from October 2014 to March 2022. Using a shorter timeframe from March 2017 to April 2020, Hao et al. [54] looked at how contagion transmits back and forth between BTC, ETH and three US market indices - S&P 500 (SPX), Dow Jones Industrial Average (DJI) and Nasdaq 100 (NDX) - using the Joe-Clayton copula GARCH model along with nonlinear Granger causality test. In another study [17], the regime switching skew-normal model was adopted to investigate the linear and non-linear contagion among big financial markets such as Euronext100, FTSE100 and Nikkei225 over a 3-year period from 2015 to 2018. Despite the differences in methodology and data, a consistent finding across these studies is that contagion intensifies during periods of heightened uncertainty, such as market crashes, major financial announcements and significant monetary policy changes.
@@
Beyond analyzing contagion effects between individual assets, researchers have also turned their attention to understanding this phenomenon at the community level. For instance, [20] investigated how different types of community structure formed from a set of banks influence the spread of liquidity contagion among the banks. The core research question tested is whether these community structures amplify or inhibit financial contagion. Using a micro-structural network model, the authors simulated how liquidity shocks propagate under various network topologies that differ in the strength of their community structure (e.g. the degree to which groups of nodes are tightly connected within their group and loosely connected to nodes outside their group). They found that the presence of communities generally increases both the number of banks affected and the overall liquidity shortfall, with contagion being most severe in networks with an intermediate level of community clustering. In this regard, the authors suggested that this result reveals a non-linear (U-shaped) relationship between community strength and contagion severity. In particular, a structure with individually strong and segregated communities can prevent shocks from spreading to the entire system, resulting in less overall contagion. However, as the community structure weakens to an intermediate level, shocks can reverberate within the remaining tighter communities, increasing their magnitude before spreading to the rest of the network, thus amplifying the initial shock. In contrast, shocks from a network with a weaker community structure dissipate rapidly without generating chain reactions. The study also tested the effectiveness of several policy interventions, including raising liquidity requirements system across (i.e. among banks across different communities) or within specific communities (i.e. among banks within a community), assigning liquidity buffers based on centrality measures, and increasing market confidence through transparency. Results showed that higher liquidity requirements and improved market trust significantly reduce contagion, especially in tightly connected communities. Overall, this study emphasizes that regulators must consider the underlying community structures when assessing systemic risk and designing regulatory policies.
@@ 2.3 Noise in Financial Assets and Noise Removal in Their Correlations
Given the existence of noise in financial assets and its distortion in financial correlations, a widely used method for noise removal relies on Random Matrix Theory (RMT) and the Marcenko-Pastur distribution [26]. This approach, introduced by Marcenko and Pastur, states that the eigenvalues of a purely random correlation matrix follow a specific distribution [66]. Therefore, according to the theory, if all eigenvalues of an empirical matrix fall within this range, the matrix contains only noise. Conversely, eigenvalues outside this range are considered to carry meaningful information. Noise is removed by eliminating the eigenvalues within the Marcenko–Pastur bounds using techniques like Linear Shrinkage, Eigenvalue Clipping, or Rotationally Invariant Estimators [67]. Indeed, this noise removal approach has been employed successfully in various studies. For instance, Laloux et al. in their portfolio optimization task denoised a correlation matrix of 406 S&P500 stocks [68] - a main variable in the Markowitz portfolio optimization algorithm, achieving better portfolio diversification with lower risk and higher returns compared to the original one. On the other hand, Laura et al. used it to identify safe-haven portfolios during market stress [4], while Hirdesh et al. found clearer correlation patterns, aiding in distinguishing different market states, such as stable and market crash [69]. Furthermore, this method has also been used in other fields such as biology [70], medicine [71], and physics [72]. However, a key limitation of this method is that it is sensitive to the length of the time period $T$ and the number of time series $N$. As a result, a change in either of them can affect the removal decision [73].
@@
To address this limitation, an alternative approach using Tracy–Widom theory has been introduced [74]. In brief, Tracy and Widom in [29] originally identified a fixed distribution for the largest eigenvalue of a random covariance matrix. Subsequent studies extended this to other eigenvalues due to their recurrent properties [75]. As a result, each eigenvalue’s distribution in a random matrix is characterized. Thus, given an empirical matrix, its noisy and informative eigenvalues can be distinguished by comparing them with corresponding random matrix-based theoretical distributions. Unlike Marcenko–Pastur, this approach is less sensitive to data size even in very small cases like $T=N=5$ and especially shows robustness when $T\geq 200$ and $N\geq 200$ [76]. Moreover, it incorporates empirical evidence, such as eigenvalue distributions obtained from thousands of simulations, into the noise removal process [75]. Therefore, it relies on observed distributions rather than strictly theoretical bounds as with the case of Marcenko–Pastur [75], making it more robust and resilient. This theory has been adopted widely in different domains. For example, it was used for fault detection in the aviation field by comparing the largest empirical eigenvalue to the Tracy–Widom theoretical distribution, as shown in [77]. In this regard, if the largest empirical eigenvalue falls within the 1% critical region of the Tracy–Widom distribution, it is considered to deviate from typical noise behavior, suggesting faults in aircraft sensor systems. Similarly, in a biology-related study [32], the authors isolated meaningful biological signals by identifying and removing noisy eigenvalues from covariance matrices built from single-cell genomic data, revealing underlying characteristics hidden by the noise. In finance, several studies have demonstrated the effectiveness and robustness of Tracy-Widom theory in identifying noisy components within covariance or correlation matrices derived from financial asset data [78, 2].

### v2

@@ Abstract
As global financial markets become increasingly interconnected, financial contagion has developed into a major influencer of asset price dynamics. Motivated by this context, our study explores financial contagion both within and between asset communities. We contribute to the literature by examining the contagion phenomenon at the community level rather than among individual assets. Our experiments rely on high-frequency data comprising cryptocurrencies, stocks and US ETFs over the 4-year period from April 2019 to May 2023. Using the Louvain community detection algorithm, Vector Autoregression contagion detection model and Tracy-Widom random matrix theory for noise removal from financial assets, we present three main findings. Firstly, while the magnitude of contagion remains relatively stable over time, contagion density (the percentage of asset pairs exhibiting contagion within a financial system) increases. This suggests that market uncertainty is better characterized by the transmission of shocks more broadly than by the strength of any single spillover. Secondly, there is no significant difference between intra- and inter-community contagion, indicating that contagion is a system-wide phenomenon rather than being confined to specific asset groups. Lastly, certain communities themselves, especially those dominated by Information Technology assets, tend to act as major contagion transmitters in the financial network over the examined period, spreading shocks with high densities to many other communities. Our findings suggest that traditional risk management strategies such as portfolio diversification through investing in low-correlated assets or different types of investment vehicle might be insufficient due to widespread contagion.
@@
[label2]organization=School of Computing, Dublin City University, addressline=Collins Ave Ext, Whitehall, city=Dublin, postcode=D09 V209, country=Ireland
@@ 1 Introduction
In light of the inherent risks associated with investing, a substantial body of literature has sought to uncover the underlying characteristics of financial markets to help investors make more informed decisions. These studies focus on enhancing investment strategies by identifying factors that can help in mitigating risks and optimizing returns [11, 12, 13]. One significant area of this topic is the investigation of financial contagion - the phenomenon where economic shocks or market disturbances in one asset, market sector or country’s financial market spread to other assets or sectors in that country or even other countries, leading to price fluctuations in the affected ones and potentially amplifying systemic risk [14]. Understanding these contagion mechanisms is crucial, as they can undermine diversification strategies and exacerbate market volatility. For example, a study conducted by Luchtenberg and Vu [15] discovered the financial contagion effects between 10 countries from North America, Europe and East Asia Pacific during one of the biggest financial crashes in the Global Financial Crisis (GFC) of 2007/08. They revealed that this phenomenon became evident in all considered countries with key economic fundamentals such as trade structure, inflation, interest rates and investor risk aversion being significant determinants of contagion. The results suggest that international diversification by geography alone may not effectively shield investors, and that policymakers in both developed and developing countries must consider contagion risks in their financial strategies. In [16], the same pattern occurred when the contagion was shown to strike rapidly during the outbreak of the Covid-19 pandemic across global financial markets, especially in developing countries, eventually causing the synchronized downturns of numerous assets across various investment vehicles. To this end, our study contributes to this line of inquiry by examining the contagion dynamics across traditional assets such as stocks and ETFs as well as emerging ones like cryptocurrencies.
@@
This study draws on high-frequency data consisting of 221 time series of closing prices recorded at 30-minute intervals from April 2019 to May 2023. This period is chosen for its inclusion of several major events - both positive and negative - that significantly impact the financial markets. Moreover, it is also a time frame analyzed in related research [12]. The dataset covers 146 stocks, 49 U.S.-based exchange-traded funds (ETFs) tied to corresponding U.S. indices and 27 cryptocurrencies. All of this data was collected from the FirstRate Data platform. Regarding the methodologies, we use Minimum Spanning Tree (MST) and Louvain community detection to identify the community structure of the assets [26, 11]. The contagion signals are detected using the well-known Vector Autoregression (VAR) approach [27]. In addition, measurement noise and inherent noise have been proven to exist in financial markets and potentially impact relevant experimental results [28, 26]. Therefore, to remove this effect from assets, we adopt the Tracy-Widom random matrix theory, which was introduced in [29] and has been widely recognized as an efficient theory for noise removal in various areas, especially financial markets [30, 31, 32].
The remainder of the article is organized as follows: Section 2 presents an overview of related works. Section 3 provides a description of the datasets. Section 4 discusses methodology, preprocessing procedures and metrics. Section 5 describes the experimental results followed by their implications and hypothesis. Section 6 provides robustness analyses. Lastly, the conclusion of this study is given in Section 7.
@@ 2.1 Upheaval in Financial Markets due to Recent Events
Global financial markets were affected by a series of major events during the 4-year period from April 2019 to May 2023. One of those that generated a great deal of headlines was the US-China trade war, which began in January 2018 with the US imposing tariffs on various Chinese imports, especially in the Information Technology sector [33]. In response, China enacted reciprocal tariffs on US goods and introduced legislation restricting the use of US-made technology products by Chinese consumers  [34, 35]. The trade war lasted for more than 2 years until January 2020, peaking in intensity during 2019 and negatively affected financial markets in both countries. Sectors such as Commodities, Energy and Technology experienced sharp declines, with major stocks and ETFs suffering substantial losses [36, 37]. Furthermore, as the central hubs of the global economy, the effects of this event spilled over to other parts of the world, increasing the volatility of financial markets worldwide [38].
The next noticeable event is the Covid-19 pandemic [39]. Different studies have explored its impact across various aspects, such as asset correlations, investor behavior, policy responses and market reactions [40, 41, 42]. The same conclusion made by the majority is that the pandemic had an exceptionally strong and widespread effect on global financial markets, leading to one of the most significant economic crises in history. Specifically, this event caused a sharp price decline in a wide range of assets across nearly all types of investment vehicle [40, 41], with the most severe effects observed during the pandemic’s first outbreak, between March and June 2020 [40, 41]. These consequences largely resulted from government interventions aimed at containing the virus, such as international border closures, mobility restrictions, and widespread shutdowns of economic activity [42].
Another event worth mentioning is the ongoing Ukraine-Russia conflict, which occurred on 24 February 2022 [43]. Russia, as a major exporter of oil and natural gas, while Ukraine, a key energy transit route and grain supplier, both play important roles in global supply chains, especially for countries in Europe. Consequently, the conflict between the two countries had a significant impact on financial markets, especially in relevant business sectors such as Energy, Financials and Consumer Staples [44, 45]. However, it is shown that the effects were uneven since the most severe financial consequences concentrated on European countries due to their strong commercial and political linkages with Russia and Ukraine[46, 47], while the impact on Asian markets and developed economies like the United States was limited or negligible [48, 49, 50]. Notably, the financial impact of this event was found to be modest compared to the past global disruptions like Covid-19 and the US–China trade war [46, 49, 51].
@@ 2.2 Contagion in Financial Markets
The topic of financial contagion has been extensively explored in the literature, with studies examining various dimensions of the phenomenon, such as contagion across countries, market indicators (e.g., interest rates, inflation rates, exchange rates, etc), trading exchanges and individual assets, both within the same investment type and across different types. For example, Ahmed and Masahiro in [52] examined financial contagion among 9 Asian countries using 3 market indicators: stock market index, interest rate, and exchange rate against the US dollar. They employed the Vector Autoregressive (VAR) model to detect possible contagion effects during the period from July 1997 to June 1998, covering the 1997 Asian Crisis. Their findings revealed contagion in stock indices and currency exchange rates. However, the effect was weak and limited since not all examined countries showed signs of receiving or transmitting contagion while those that did only interacted with a few others. By contrast, they found no evidence of contagion in interest rates. Interestingly, the study noted that while contagion intensified during periods of high market turbulence, it was not the primary driver of the crisis in 1997 due to its weak and limited effect seen among the countries.
@@
As a common method for measuring contagion, Tugba et al. [53] also used the VAR model for their research. In this study, they investigated the shock spillovers between cryptocurrencies, technology stocks, US ETFs and NFTs from March 2018 to September 2021. Their study assessed contagion both within and across different types of investment vehicle. The results showed that contagion primarily occurs within the same investment type since most contagion links are found among assets of the same type. However, these contagions do appear between technology stocks and US ETFs, while very little spillover is found between cryptocurrencies and traditional assets. In addition, NFTs show no contagion with other investment types. Focusing on a different aspect of contagion, Giudici and Pagnottoni in [54] calculated the total spillover index (TSI) and directional spillover indices (DSI) using the Kwiatkowski–Phillips–Schmidt–Shin (KPPS) forecast error variance decompositions to examine hourly Bitcoin return spillovers among 5 cryptocurrency exchanges between July 2017 and June 2018. They emphasized that the spillovers among these exchanges constantly change over time. Remarkably, Bitfinex and Gemini exchanges are identified as leaders in price formation, transmitting return spillovers to the rest, while Bittrex tends to be a follower, receiving shocks from the others.
@@
Apart from VAR, different methods have been introduced to examine financial contagion across different time periods and market conditions. For instance, Ahelegbey et al. introduced a new variant of the VAR model, known as Network VAR, in which the incorporation of interconnectedness networks into the original framework helps filter out unreliable contagion effects between time series [55]. Using this novel approach, they investigated the contagions among 20 major stock markets worldwide and further examined how these interconnections relate to global market risks over the period from 1999 to 2021. In a related study, the authors of [19] employed DCC-GARCH (Dynamic Conditional Correlation GARCH) and Wavelet Cross-Correlation methods to examine the contagion across various cryptocurrencies and equity markets in both developing and giant economies from October 2014 to March 2022. Using a shorter timeframe from March 2017 to April 2020, Hao et al. [56] looked at how contagion transmits back and forth between BTC, ETH and three US market indices - S&P 500 (SPX), Dow Jones Industrial Average (DJI) and Nasdaq 100 (NDX) - using the Joe-Clayton copula GARCH model along with nonlinear Granger causality test. In another study [17], the regime switching skew-normal model was adopted to investigate the linear and non-linear contagion among big financial markets such as Euronext100, FTSE100 and Nikkei225 over a 3-year period from 2015 to 2018. Despite the differences in methodology and data, a consistent finding across these studies is that contagion intensifies during periods of heightened uncertainty, such as market crashes, major financial announcements and significant monetary policy changes.
@@
Beyond analyzing contagion effects between individual assets, researchers have also turned their attention to understanding this phenomenon at the community level. For instance, [20] investigated how different types of community structure formed from a set of banks influence the spread of liquidity contagion among the banks. The core research question tested is whether these community structures amplify or inhibit financial contagion. Using a micro-structural network model, the authors simulated how liquidity shocks propagate under various network topologies that differ in the strength of their community structure (e.g. the degree to which groups of nodes are tightly connected within their group and loosely connected to nodes outside their group). They found that the presence of communities generally increases both the number of banks affected and the overall liquidity shortfall, with contagion being most severe in networks with an intermediate level of community clustering. In this regard, the authors suggested that this result reveals a non-linear (U-shaped) relationship between community strength and contagion severity. In particular, a structure with individually strong and segregated communities can prevent shocks from spreading to the entire system, resulting in less overall contagion. However, as the community structure weakens to an intermediate level, shocks can reverberate within the remaining tighter communities, increasing their magnitude before spreading to the rest of the network, thus amplifying the initial shock. In contrast, shocks from a network with a weaker community structure dissipate rapidly without generating chain reactions. The study also tested the effectiveness of several policy interventions, including raising liquidity requirements across (i.e. among banks across different communities) or within specific communities (i.e. among banks within a community), assigning liquidity buffers based on centrality measures, and increasing market confidence through transparency. Results showed that higher liquidity requirements and improved market trust significantly reduce contagion, especially in tightly connected communities. Overall, this study emphasizes that regulators must consider the underlying community structures when assessing systemic risk and designing regulatory policies.
@@ 2.3 Noise in Financial Assets and Noise Removal in Their Correlations
Given the existence of noise in financial assets and its distortion in financial correlations, a widely used method for noise removal relies on Random Matrix Theory (RMT) and the Marchenko-Pastur distribution [26, 68]. This approach, introduced by Marchenko and Pastur, states that the eigenvalues of a purely random correlation matrix follow a specific distribution [69]. Therefore, according to the theory, if all eigenvalues of an empirical matrix fall within this range, the matrix contains only noise. Conversely, eigenvalues outside this range are considered to carry meaningful information. Noise is removed by eliminating the eigenvalues within the Marchenko–Pastur bounds using techniques like Linear Shrinkage, Eigenvalue Clipping, or Rotationally Invariant Estimators [70]. Indeed, this noise removal approach has been employed successfully in various studies. For instance, Laloux et al. in their portfolio optimization task denoised a correlation matrix of 406 S&P500 stocks [71] - a main variable in the Markowitz portfolio optimization algorithm, achieving better portfolio diversification with lower risk and higher returns compared to the original one. On the other hand, Laura et al. used it to identify safe-haven portfolios during market stress [4], while Hirdesh et al. found clearer correlation patterns, aiding in distinguishing different market states, such as stable and market crash [72]. Furthermore, this method has also been used in other fields such as biology [73], medicine [74], and physics [75]. However, a key limitation of this method is that it is sensitive to the length of the time period $T$ and the number of time series $N$. As a result, a change in either of them can affect the removal decision [76].
@@
To address this limitation, an alternative approach using Tracy–Widom theory has been introduced [77]. In brief, Tracy and Widom in [29] originally identified a fixed distribution for the largest eigenvalue of a random covariance matrix. Subsequent studies extended this to other eigenvalues due to their recurrent properties [78]. As a result, each eigenvalue’s distribution in a random matrix is characterized. Thus, given an empirical matrix, its noisy and informative eigenvalues can be distinguished by comparing them with corresponding random matrix-based theoretical distributions. Unlike Marchenko–Pastur, this approach is less sensitive to data size even in very small cases like $T=N=5$ and especially shows robustness when $T\geq 200$ and $N\geq 200$ [79]. Moreover, it incorporates empirical evidence, such as eigenvalue distributions obtained from thousands of simulations, into the noise removal process [78]. Therefore, it relies on observed distributions rather than strictly theoretical bounds as with the case of Marchenko–Pastur [78], making it more robust and resilient. This theory has been adopted widely in different domains. For example, it was used for fault detection in the aviation field by comparing the largest empirical eigenvalue to the Tracy–Widom theoretical distribution, as shown in [80]. In this regard, if the largest empirical eigenvalue falls within the 1% critical region of the Tracy–Widom distribution, it is considered to deviate from typical noise behavior, suggesting faults in aircraft sensor systems. Similarly, in a biology-related study [32], the authors isolated meaningful biological signals by identifying and removing noisy eigenvalues from covariance matrices built from single-cell genomic data, revealing underlying characteristics hidden by the noise. In finance, several studies have demonstrated the effectiveness and robustness of Tracy-Widom theory in identifying noisy components within covariance or correlation matrices derived from financial asset data [81, 2].